A chemist mixes 71.1 g of carbon tetrachloride with 65.4 g of isopropenylbenzene and 4.02 g of 2-ethyltoluene. Calculate the percent by mass of each component of this solution. Be sure each of your answer entries has the correct number of significant digits mass percentー1 □ component 0% carbon tetrachloride isopropenylbenzene 2-ethyltoluene




1
Expert's answer
2021-06-01T02:45:15-0400

Solution:

Mixture mass = Mass of carbon tetrachloride + Mass of isopropenylbenzene + Mass of 2-ethyltoluene

Mixture mass = 71.1 g + 65.4 g + 4.02 g = 140.52 g


%carbon tetrachloride = (Mass of carbon tetrachloride / Mixture mass) × 100%

%carbon tetrachloride = (71.1 g / 140.52 g) × 100% = 50.6%

%isopropenylbenzene = (Mass of isopropenylbenzene / Mixture mass) × 100%

%isopropenylbenzene = (65.4 g / 140.52 g) × 100% = 46.5%

%2-ethyltoluene = (Mass of 2-ethyltoluene / Mixture mass) × 100%

%2-ethyltoluene = (4.02 g / 140.52 g) × 100% = 2.9%

50.6% + 46.5% + 2.9% = 100%


Answer:

%carbon tetrachloride = 50.6%

%isopropenylbenzene = 46.5%

%2-ethyltoluene = 2.9%